# Qudit inspired optimization for graph coloring

### v1

@@ IV Conclusion
We compared our results for a range of GCPs to some of the state-of-the-art methods in the literature schuetz_22_2 ; wang_23 , and both QdLQA and QdGD find equally good or better solutions for all of the test cases, with the exception of the Pubmed graph. We also tested the algorithms on graphs with large chromatic numbers hebrard_19 . Even though neither algorithm solved these problems completely, they still provided reasonable quality solutions (with normalized error $\epsilon\approx 10^{-3}$) using modest computing resources such as a desktop or a laptop. Notably, QdLQA almost always provided equally good or better solutions than QdGD (with the exception of the Cora and the Facebook graph). However, QdGD was heuristically found to be more efficient, since it minimizes the desired cost function directly. We refer the reader to Appendix A for further details.
There are many compelling possible extensions to our work. For example, each local Hilbert space dimension can be set manually, and thus, the approach can easily be extended to integer optimization problems requiring different numbers of states on each node. Furthermore, the interpretation of a vector on a sphere might provide new insights into why certain solutions are preferred (which can also be done with an analysis of the solutions and the interactions due to $h$). Additionally, there is also plenty of room for general optimizations, hyperparameter tuning, and different choices of, e.g., initial configurations. It will also be important to investigate if the algorithms can provide competitive solutions to GCPs containing millions of nodes, and see if they can be generalized to other problems, like large-scale graph-based machine learning tasks zheng2020distdgl ; hu2020open .
@@ Acknowledgment
Appendix A Computer architectures and runtimes
@@
All computations, except those on the PubMed, Facebook, and Wiki-vote graphs, were run on a laptop with an 11th Gen Intel(R) Core(TM) i5-1135G7 @ 2.40GHz processor, whereas the latter ran on a desktop with an 11th Gen Intel(R) Core(TM) i5-11500 @ 2.70GHz processor. In Fig. 5, we plot the average computation time for one run for the different graphs for the parameters used in this work, see Appendix B. Note that parameters like $N_{\textrm{runs}}$ and $\alpha$ vary, so the figure is an illustration of the cost of obtaining the presented data, rather than fairly comparing the computational cost for the different graphs. Figure 5(a) shows the average time for one run of the algorithms for the different graphs on the laptop and Figure 5(b) on a stationary computer  [^1] The partially large variances for QdLQA (and possibly for QdGD) comes from the fact that other processes were run simultaneously to the algorithms on the corresponding hardware.. As expected, the run time is significantly impacted by the Patience hyperparameter for QdGD, which allows the algorithm to stop early.
@@
Figure 5: Average computation times on a (a) laptop (b) stationary computer for $N_{\textrm{runs}}$ of QdLQA and QdGD for different graphs. The error bars correspond to the standard deviation of the data.
Appendix B Hyperparamters
There are several hyperparameters needed in both the QdLQA and the QdGD algorithms. These and their effects are summarized in Table 3. Additionally, the Adam optimizer has further hyperparameters that could be tuned (here, we only adjust the learning rate), however, in this work, they are left to their default values in PyTorch.
In Table 4, we show the hyperparameters used to produce the data in the Tables 1 and 2. In most cases, we found our default values $\eta=0.5$, $h=3.0$, $N_{\textrm{Steps}}=1000$ and $\gamma=1.0$ work well, however, for some graphs, we either started with the default values and tuned them manually or did a grid search. Note that $N_{\textrm{steps}}=3175$ for QdGD for the queen13-13 graph was chosen such that we get equally many gradient descent steps as in QdLQA with $N_{\textrm{steps}}=1000$ and $\alpha=2$.
@@ Appendix B Hyperparamters
$N_{steps}$ QdLQA Number of steps used when going from $t=0$ to $t=1$ in the quantum adiabatic evolution.
$N_{steps}$ QdGD Number of gradient descent steps.
$\gamma$ Both Prefactor in $E_{\textrm{W}}(\theta)$ from Eq. (10).
$\alpha$ QdLQA Determines the number of gradient descent steps at each time step $t$ as $round(e^{t\alpha})$.
$\eta$ Both Learning rate in the Adam optimizer.
$f$ QdLQA Determines the perturbation in the angles in the initial state.
$h$ Both Determines the perturbation in the interactions in the Potts model, see Eq. (9).
$N_{runs}$ Both Number of times we run the algorithm.
Patience QdGD If no improvement in $E_{\textrm{Potts}}$ is seen after this many gradient descent steps, the algorithm stops.
$\tilde{f}$ QdGD Determines the sampling interval for each qudit in the intial state.
@@
Graph $N_{steps}$ $\gamma$ $\alpha$ $\eta$ $f$ $h$ $N_{runs}$
myciel5 1000 1.0 0.0 0.5 0.0 3.0 100
myciel6 1000 1.0 0.0 0.5 0.0 3.0 100
queen5-5 1000 1.0 0.0 0.5 0.0 3.0 100
queen6-6 1000 1.0 0.0 0.5 0.0 3.0 100
queen7-7 1000 1.0 0.0 0.5 0.0 3.0 100
queen8-8 1000 1.0 0.0 0.5 0.1 3.0 100
queen9-9 1000 1.0 0.0 0.5 0.0 3.0 100
queen8-12 1000 1.0 0.0 0.5 0.0 3.0 100
queen11-11 1000 1.0 0.0 0.5 0.1 3.0 100
queen13-13 1000 0.75 2.0 0.5 0.0 3.0 100
Cora 1000 0.25 0.0 0.5 0.1 10.0 100
Pubmed 1000 0.1 0.0 0.1 0.5 3.0 100
email-Eu-core 1500 0.5 0.0 0.1 0.1 3.0 100
Facebook 1000 0.1 0.0 0.1 0.1 3.0 100
Wiki-vote 1000 0.1 0.0 0.5 0.1 3.0 100
@@
$N_{steps}$ $\gamma$ Patience $\eta$ $\tilde{f}$ $h$ $N_{runs}$
@@
Table 4: Parameters used to compute the data in Tables 1 and 2.

### v2

@@ IV Conclusion
We compared our results for a range of GCPs to some of the state-of-the-art methods in the literature schuetz_22_2 ; wang_23 , and both QdLQA and QdGD find equally good or better solutions for all of the test cases, with the exception of the PubMed graph. We also tested the algorithms on graphs with large chromatic numbers hebrard_19 . Even though neither algorithm solved these problems completely, they still provided reasonable-quality solutions (with normalized error $\epsilon\approx 10^{-3}$) using modest computing resources such as a desktop or a laptop. Notably, QdLQA almost always provided equally good or better solutions than QdGD (with the exception of the Cora and the Facebook graph). However, QdGD was heuristically found to be more efficient, since it minimizes the desired cost function directly. We refer the reader to Appendix B for further details.
There are many compelling possible extensions to our work. For example, each local Hilbert space dimension can be set manually, and thus, the approach can easily be extended to integer optimization problems requiring different numbers of states on each node. It general, it would also be interesting to do a detailed comparison between local quantum annealing and simulated (thermal) annealing; however, it seems that for the GCP, the latter suffers convergence and scalability issues kose2017simulated ; nolte1996simulated ; lukasik2008parallel . Furthermore, the interpretation of a vector on a sphere might provide new insights into why certain solutions are preferred (which can also be done with an analysis of the solutions and the interactions due to $h$). In addition, there is plenty of room for general optimizations, hyperparameter tuning, and different choices (e.g., initial configurations). We have chosen to initialize the nodes in the ground states of $-\hat{L}^{x}_{i}$ since this connects with quantum annealing for qudit systems. However, other initial states, e.g., all colors equally populated, might improve the results further. It will also be important to investigate if the algorithms can provide competitive solutions to GCPs containing millions of nodes and see if they can be generalized to other problems, such as large-scale graph-based machine learning tasks zheng2020distdgl ; hu2020open .
Finally, we note that by restricting the cost function to be a quantum Hamiltonian (see, for example, Ref. deller_23 ), our method is similar to simulating quantum annealing in a qudit system (without entanglement). Therefore, one could also investigate if allowing for entanglement in the system might help the algorithms find better solutions. One approach would be to use tensor networks with a bond dimension larger than 1 (see, for example, Ref. bauer_15 ). However, this comes with significant computational overhead and might be unfeasible to scale to industrial-size problems.
Data availability: The data supporting our results can be found at https://zenodo.org/records/11483787.
@@ Acknowledgment
Appendix A Pseudocode
Here, we present pseudocode for the algorithms QdLQA and QdGD (Algorithm 1 and Algorithm 2, respectively). In both algorithms, we first load the graph’s nodes and edges and initialize the hyperparameters and an empty list $e_{\textrm{list}}$. For each of the $N_{\textrm{runs}}$ of the algorithms, we initialize the state $\psi$ which depends on the parameters $\vec{\phi}$. In practice, this can, for example, be a list of length |$\absolutevalue{V}$ | (the number of nodes), where each entry is a vector and depends on elements of $\vec{\phi}$ as written in Eq. (6). In this work, the vectors are initialized to the ground state of $-\hat{L}^{x}_{i}$ for QdLQA and random values for QdLQA. However, other initialization schemes are possible. For QdLQA, we then let the time $t$ go from 00 to $1$, and at each time step, we minimize a cost function $E_{\textrm{Total}}(\psi,t)$ using $\alpha$ gradient descent steps. We then compute the classical energy $e_{\textrm{Potts}}$ (we use-upper case letters for functions and lower-case letters for variables). At the end of each run, the lowest energy is added to $e_{\textrm{list}}$, and finally, $e_{\textrm{list}}$ is returned. This can then be used to extract the values in Table 1 and Table 2. For QdGD, we proceed in a similar manner. Here, however, the cost function does not depend on time, and we also stop the algorithm if no improvement in classical energy is found based on the hyperparameter Patience.
Load $\alpha$, $f$, $N_{\text{runs}}$, $N_{\text{steps}}$
$e_{\textrm{list}}$ is an empty list
for $i=1$ to $N_{\text{runs}}$ do
      Initialize the state $\psi$ that depend on parameters $\vec{\phi}$ with perturbation $f$
      $t\leftarrow 0$
      $e_{\text{best}}\leftarrow 1000000$
      while t < 1 do
            Minimize $E_{\textrm{Total}}(\psi,t)$ with $\alpha$ steps
            $e_{\text{Potts}}\leftarrow E_{\text{Potts}}(\psi)$
            $e_{\text{best}}\leftarrow\text{min}(e_{\text{best}},e_{\text{Potts}})$
            if $e_{\text{best}}$ = 0 then
                  break
            end if
           $t\leftarrow t+1/N_{\text{steps}}$
      end while
     Add $e_{\text{best}}$ to $e_{\textrm{list}}$
end for
Return $e_{\textrm{list}}$
For the specifics regarding the initialization of the state $\psi$ and formulation of the cost function $E_{\textrm{Total}}(\psi,t)$, see the main text.
Algorithm 1 QdLQA
For the specifics regarding the initialization of the state $\psi$ and formulation of the cost function $E_{\textrm{Total}}(\psi,t)$, see the main text.
Initialize $\tilde{f}$, $N_{\text{runs}}$, $N_{\text{steps}}$, Patience
$e_{\textrm{list}}$ is an empty list
for $i=1$ to $N_{\text{runs}}$ do
      Initialize the state $\psi$ that depend on parameters $\vec{\phi}$ with perturbation $\tilde{f}$
      $e_{\text{best}}\leftarrow 1000000$
      for $j=1$ to $N_{\text{steps}}$ do
            Minimize $E_{\textrm{Total}}(\psi)$ with one step
            $e_{\text{Potts}}\leftarrow E_{\text{Potts}}(\psi)$
            $e_{\text{best}}\leftarrow\text{min}(e_{\text{best}},e_{\text{Potts}})$
            if $e_{\text{best}}$ = 0 or no improvement  then
                  break
            end if
      end for
     Add $e_{\text{best}}$ to $e_{\textrm{list}}$
end for
Return $e_{\textrm{list}}$
For the specifics regarding the initialization of the state $\psi$ and formulation of the cost function $E_{\textrm{Total}}(\psi)$, see the main text.
Algorithm 2 QdGD
For the specifics regarding the initialization of the state $\psi$ and formulation of the cost function $E_{\textrm{Total}}(\psi)$, see the main text.
The pseudocode for the cost function we use for QdLQA is shown in Algorithm 3 (for QdGD it is just the same function evaluated at $t=1$). $\hat{L}^{x}$ is a matrix of size $c\times c$ and is given by the $x$ angular momentum operator (we drop the subindex $i$ since it is site independent in our implementation), see Sec. II.2, and $j_{\textrm{max}}$ is a fixed qudit initialized to one color. Note that the $j_{\textrm{max}}$ qudit can also be emitted in the second for loop as it only contributes with a constant term. In the algorithm, squaring and taking the logarithm are done componentwise. Lastly, we show how we extract the final color configuration in Algorithm 4. There, $conf$ is the zero vector of size |$\absolutevalue{V}$ | and argmax gives the index of the component with the largest magnitude of the vector $\psi[i]^{2}$.
Load vertices $V$, edges $\mathcal{E}$, matrix $\hat{L}^{x}$, $j_{\textrm{max}}$, $\gamma$, $h$
$e_{\textrm{F}}\leftarrow 0$
$e_{\textrm{W}}\leftarrow 0$
$e_{\textrm{I}}\leftarrow 0$
for $(i,j)\in\mathcal{E}$ do
[…]
      $e_{\textrm{F}}\leftarrow e_{\textrm{F}}+\psi[i]^{2}\cdot\psi[j]^{2}(\mathcal{E}[i,j]+h_{i,j})$
end for
for $l=1$ to |$\absolutevalue{V}$ | do
      $e_{\textrm{W}}\leftarrow e_{\textrm{W}}+\gamma\psi[l]^{2}\cdot\log(\psi[l]^{2})$
end for
for $n=1$ to |$\absolutevalue{V}$ | do
      if $n\neq j_{\textrm{max}}$  then
            $e_{\textrm{I}}\leftarrow e_{\textrm{I}}-\psi[n]^{T}\cdot L^{x}\cdot\psi[n]$
      end if
end for
Return $(1-t)e_{\textrm{I}}+t(e_{\textrm{F}}+e_{\textrm{W}})$
Computing the cost function from Eq. (11). $t$ is a floating point and $\psi$ is, for example, a list of vectors and depends on the parameters $\vec{\phi}$ so that the $i$-th entry is given by Eq. (6).
Algorithm 3 $E_{\textrm{Total}}(t,\psi)$
Computing the cost function from Eq. (11). $t$ is a floating point and $\psi$ is, for example, a list of vectors and depends on the parameters $\vec{\phi}$ so that the $i$-th entry is given by Eq. (6).
Load vertices $V$, edges $\mathcal{E}$
$conf\leftarrow\vec{0}$
for $i=1$ to |$\absolutevalue{V}$ | do
      $conf[i]\leftarrow\textrm{argmax}(\psi[i]^{2})$
end for
$e_{\textrm{Potts}}\leftarrow 0$
for $(i,j)\in\mathcal{E}$ do
      if $conf[i]=conf[j]$  then
            $e_{\textrm{Potts}}\leftarrow e_{\textrm{Potts}}+1$
      end if
end for
Return $e_{\textrm{Potts}}$
Computing the energy from Eq. (1). $\psi$ is, for example, a list of vectors and depends on the parameters $\vec{\phi}$ so that the $i$-th entry is given by Eq. (6).
Algorithm 4 $E_{\textrm{Potts}}(\psi)$
Computing the energy from Eq. (1). $\psi$ is, for example, a list of vectors and depends on the parameters $\vec{\phi}$ so that the $i$-th entry is given by Eq. (6).
Appendix B Computer architectures and runtimes
@@
All computations, except those on the PubMed, Facebook, and Wiki-vote graphs, were run on a laptop with an 11th-generation Intel(R) Core(TM) i5-1135G7 @ 2.40GHz processor, whereas the latter ran on a desktop with an 11th-generation Intel(R) Core(TM) i5-11500 @ 2.70GHz processor. In Fig. 6, we plot the average computation time for one run for the different graphs for the parameters used in this work; see Appendix C. Note that the parameter $\alpha$ varies (for queen13-13), so the figure is an illustration of the cost of obtaining the data presented, rather than fairly comparing the computational cost for the different graphs. Figure 6(a) shows the average time for one run of the algorithms for the different graphs on the laptop and Figure 6(b) on a stationary computer ***The partially large variances for QdLQA (and possibly for QdGD) comes from the fact that other processes were run simultaneously to the algorithms on the corresponding hardware.. As expected, the run time is significantly impacted by the Patience hyperparameter for QdGD, which allows the algorithm to stop early.
For a more thorough comparison between the graph neural network approaches, QdLQA, and QdGD, we run the PI-SAGE algorithm from Ref. schuetz_22_2 on the previously mentioned laptop for several datasets. For the calculations, we use the code available at gnn_git and the parameters reported in Ref. schuetz_22_2 . In Fig. 7, we show the average runtimes; as the problem sizes increase, the QdLQA and QdGD run significantly faster (note that this might change on different computer architectures, where, for example, GPUs are available). In addition, the QdLQA and QdGD only optimize (|$(\absolutevalue{V}-1)(c-1)$ parameters (the minus one in (|$(\absolutevalue{V}-1)$ comes from already assigning one color to one qudit in the initial state), which is a fraction of what, for example, PI-SAGE requires (even though the number of parameters in principle scales as $\sim c+\absolutevalue{V}$ |). For example, for the queen5-5, queen9-9, and queen11-11 datasets, our methods have 96, 720, and 1200 parameters, whereas the PI-SAGE method uses 7210, 12663, and 13411 parameters (using the hyperparameters from Ref. schuetz_22_2 ).
Lastly, we emphasize that the graph neural network approaches also might need $N_{\textrm{runs}}>1$ to find an optimal solution. By introducing $p^{\textrm{PI-SAGE}}_{\textrm{min}}=N_{\textrm{min}}/N_{\textrm{runs}}$, we get the following tuples (dataset, $E^{\textrm{PI-SAGE}}_{\textrm{min}}$, $p^{\textrm{PI-SAGE}}_{\textrm{min}}$): (myciel5, 0, 100/100), (myciel6, 0, 100/100), (queen5-5, 0, 85/100), (queen6-6, 0, 17/100), (queen7-7, 0, 1/100), (queen8-8, 1, 3/100), (queen9-9, 1, 6/100), (queen8-12, 2, 6/100), and (queen11-11, 13, 1/100) in our calculations (the last result actually improves on what was reported in Ref. schuetz_22_2 ).
@@
Figure 6: Average computation times on (a) a laptop and (b) a stationary computer for $N_{\textrm{runs}}$ of QdLQA and QdGD for different graphs. The error bars correspond to the standard deviation of the data.
Figure 7: Average computation times on the laptop described in the main text for QdLQA, QdGD, and PI-SAGE. For all algorithms, we use $N_{\textrm{runs}}=100$.
Appendix C Hyperparamters
@@
$N_{\text{steps}}$ QdLQA Number of steps used when going from $t=0$ to $t=1$ in the quantum adiabatic evolution
$N_{\text{steps}}$ QdGD Number of gradient descent steps
$\gamma$ Both Prefactor in $E_{\textrm{W}}(\psi)$ from Eq. (10)
$\alpha$ QdLQA Determines the number of gradient descent steps at each time step $t$
$\eta$ Both Learning rate in the Adam optimizer
$f$ QdLQA Determines the perturbation in the angles in the initial state
$h$ Both Determines the perturbation in the interactions in the Potts model, see Eq. (9)
$N_{\text{runs}}$ Both Number of times we run the algorithm
Patience QdGD If no improvement in $E_{\textrm{Potts}}$ is seen after this many gradient descent steps, the algorithm stops
$\tilde{f}$ QdGD Determines the sampling interval for each qudit in the initial state
@@
Graph $N_{\text{steps}}$ $\gamma$ $\alpha$ $\eta$ $f$ $h$ $N_{\text{runs}}$
myciel5 1000 1.0 1 0.5 0.0 3.0 100
myciel6 1000 1.0 1 0.5 0.0 3.0 100
queen5-5 1000 1.0 1 0.5 0.0 3.0 100
queen6-6 1000 1.0 1 0.5 0.0 3.0 100
queen7-7 1000 1.0 1 0.5 0.0 3.0 100
queen8-8 1000 1.0 1 0.5 0.1 3.0 100
queen9-9 1000 1.0 1 0.5 0.0 3.0 100
queen8-12 1000 1.0 1 0.5 0.0 3.0 100
queen11-11 1000 1.0 1 0.5 0.1 3.0 100
queen13-13 1000 0.75 * 0.5 0.0 3.0 100
Cora 1000 0.25 1 0.5 0.1 10.0 100
Pubmed 1000 0.1 1 0.1 0.5 3.0 100
email-Eu-core 1500 0.5 1 0.1 0.1 3.0 100
Facebook 1000 0.1 1 0.1 0.1 3.0 100
Wiki-vote 1000 0.1 1 0.5 0.1 3.0 100
@@
$N_{\text{steps}}$ $\gamma$ Patience $\eta$ $\tilde{f}$ $h$ $N_{\text{runs}}$
@@
Table 4: Hyperparameters used to compute the data in Tables 1 and 2. * For queen13-13, we conduct $\textrm{round}(e^{2t})$ gradient descent steps ($\textrm{round}()$ indicates that we round to the nearest integer) at each point in time (maximum 7).
There are several hyperparameters needed in both the QdLQA and the QdGD algorithms. These and their effects are summarized in Table 3. In addition, the Adam optimizer has further hyperparameters that could be tuned (here, we only adjust the learning rate), however, in this work, they are left to their default values in PyTorch.
In Table 4, we show the hyperparameters used to produce the data in Tables 1 and 2. Initially, we choose the default values $\eta=0.5$, $h=3.0$, $\alpha=1$, $N_{\textrm{steps}}=1000$ and $\gamma=1.0$ which work well for many of the graphs. When the algorithms neither solved the problem nor achieved better solutions than the graph neural network approaches, we either tuned the parameters manually or did a grid search. Note that for queen13-13, we choose $N_{\textrm{steps}}=3175$ for QdGD so that we get equally as many gradient descent steps as in QdLQA.